# Magnetic Localization for In-body Nano-communication Medical Systems

### v1

@@ IV-C Multilateration with Many Wires
A more robust algorithm requires a larger number of reference wires. In the following section, we present an approach feasible when we have at least nine wires. Let us consider that all the wires are parallel to one of the axes of the coordinates system and call them , , or wires, according to the axis of their parallelism.
Now, let us consider a set of parallel wires, e.g., wires, like in Fig. 2. Again, the distances calculated in the ranging phase create cylinders of possible nano-node positions along each wire. Thus, localization with Z wires enables to calculation of the (,) coordinates of the nano-node with the following set of equations:
@@
, - coordinates of the nano-node estimated with Z wires
- distance from the nano-node to the wire
@@
Then, this matrix equation can be solved using the standard least-squares approach, obtaining the estimate vector containing the and coordinates of the nano-node:
Now, the analogous approach can be performed for and wires. Having at least three wires and three wires, we can follow the algorithm given in (4)-(10) to obtain (,) and (,) coordinates. Each set of wires enables to calculation of two of the nano-node coordinates only. However, assuming the number of , , and wires is respectively equal to , , and , the final nano-node coordinates can be calculated as respective weighted means:
@@ V Simulation Results
To validate the proposed localization algorithms, a computer simulator is written in Python 3 programming language. The simulator reflects the geometric setting of DC wires being the magnetic sources (Sec. III-A) and nano-nodes equipped with magnetometers operating inside of a human body (Sec. III-B). It also takes into account the Earth’s magnetic field which might be a source of measurement errors (Sec. V-A). For each position of the nano-node inside of the body, the simulator calculates the measured magnetic field coming from each DC wire influenced by the Earth’s magnetic field. With these measurements, the position of the nano-node is estimated based on the proposed localization algorithm (Sec. IV). Finally, the localization error is calculated in comparison to the real position of the nano-node. Discussing it in more detail, each simulation follows the subsequent steps:
@@
Geometry definition. The general geometry of the simulation is created: the position of the DC wires and the relative inclination of the Earth’s magnetic field are established. While the Biot-Savart law (1) is defined for infinitely long DC wires, here, for the setup simplicity, we assume the wires are 1-2 m long (see Fig. 5-6), but their length may be increased in practical scenarios.
@@
Nano-nodes positions and inclinations. A voxel model of a 175 cm human male is implemented to emulate the human body. The model is depicted in Fig. 3; it is composed of 17,381 points in total. Each point represents a potential nano-node position for which magnetic field measurements are performed and localization is attempted. At each point, a nano-node is generated with a random orientation, which is repeated 100 times.
@@
Localization and its error. The proposed localization algorithm is executed and the position of the nano-node is estimated. The Euclidean distance between the real and estimated positions is calculated as the position error. It is averaged over 100 random nano-node orientations at each point of the human body model.
V-A Earth’s Magnetic Field
@@
A couple of models allow for the estimation of the geomagnetic field. In this paper, we use the commonly known World Mathematical Model (WMM) [43]. It was created by the U.S. National Oceanic and Atmospheric Administration’s National Centers for Environmental Information (NOAA/NCEI) and the British Geological Survey (BGS). The current model version was parameterized in 2020 and is supposed to be valid until 2025. The main goal of the WMM is to represent the planet’s magnetic field for all locations of the globe. What is quite important here, it also provides data on how large the model error is compared with the real Earth’s magnetic field. These errors for three orthogonal components (northern, eastern, and vertical ones) of the magnetic flux density are shown in Tab. II for the 2020 and 2025 years.
For the computer simulator, we calculate the average of the model error for 2020 and 2025 (see the third column of Tab. II). Then, for each nano-node, we randomly generate the WMM errors for all three components as uniformly distributed values from the range -error, +error.
@@ V-B Results
Two primary wire arrangements were investigated during simulations: with three and 15 DC wires. The electrical current for all the wires was 100 A. The created magnetic flux density was never higher than 120 mT inside the human body, as it was the maximum value in the measurement range of the used magnetometers.
The scenario with three wires was chosen, as it is the minimal number of wires when the localization is feasible; the respective algorithm is given in Sec. IV-B. On the other hand, with 15 wires, the more advanced algorithm, based on multilateration, can be verified (Sec. IV-C). This wire arrangement creates a cage-like shape where a patient can stand inside (see Fig. 4) for medical investigation. For each simulation and each nano-node, 100 simulations were performed, and the results were averaged. The localization errors are summarized in Tab. III.
In the three wires scenario, the mean position error equals 1.92 cm with a standard deviation of 0.9 cm. The detailed error map is depicted in Fig. 5. The lowest error is in the center of the phantom. The nano-nodes positioned close to the , , or planes may experience large position errors (see the human model left hand and feet in Fig. 5). It is due to the effect described in Sec. IV-B: the ranging errors may result in negative root values when calculating nano-node positions. While it might be corrected in the algorithm, the position error is still high.
The scenario with 15 DC wires produces much better position estimates; the mean error is only 0.74 cm with a standard deviation of 0.18 cm. With the error map (see Fig. 6), it can be seen that again the lowest error is in the center of the human body, but we do not have large errors on feet or hands, as the effect of negative root values does not take place here. While the mean localization accuracy (below 1 cm) might be very attractive for many medical applications, the simpler scenario with just three DC wires might be found more feasible in practice, while being also quite precise.
| Scenario | 3 wires | 15 wires |
| Errors [cm]: | ||
| Maximum X error | 6.81 | 0.850 |
| Maximum Y error | 3.33 | 0.91 |
| Maximum Z error | 6.35 | 0.59 |
| Maximum position error | 6.87 | 1.47 |
| Mean X error | 1.09 | 0.38 |
| Mean Y error | 1.13 | 0.42 |
| Mean Z error | 0.58 | 0.29 |
| Mean position error | 1.92 | 0.74 |
| Std of position error | 0.9 | 0.18 |
@@ VI Conclusions
In this paper, we presented a new magnetic-field-based localization system suitable for in-body nano-devices. In the system, external electric wires with constant current generate a magnetic field that can be measured by tiny magnetometers mounted on board the nano-devices. These measurements are sent out of the body through a body-area network and are used to calculate the nano-device positions. We proposed two versions of the localization algorithm: for three, and nine or more electric wires. We also explained how to take into account Earth’s magnetic field. We performed computer simulations for both algorithm versions showing that even with errors induced by the magnetometers and Earth’s magnetic field, the obtained localization accuracy can be about 1 cm or better. The proposed solution is a necessary step in a variety of potential applications in future medical diagnostic systems.
Future works extending these studies may go in many different directions ranging from specific system applications through nano-device hardware fine-tuning to new networking architectures. One particularly interesting example is substituting the external DC wires, generating the magnetic field, with in-body magnets. Such magnets are already considered for guiding some nano-particles [44]. The advantage of this solution is the fact that no external reference wires are required, as the source of the magnetic field is inside the body. It is, however, still to be investigated if the magnets’ positions could be sufficiently stable and the fields suitably regular and strong for accurate localization. Another related research line is the study of time characteristics of the magnetic fields generated by the external DC wires. As the static field is required for the proposed localization algorithm, it is worth examining how quickly the wires can be switched on and off, as it affects the frequency of localization measurements. These examples of future studies just indicate a wide range of research topics related to magnetic field applications for in-body medical diagnosis.
@@ Acknowledgement
This work was carried out in the framework of COST Action CA20120 INTERACT. It was also supported by the grant CBET-2039189 of US National Science Foundation and by the Polish Ministry of Science and Higher Education with the subvention funds of the Faculty of Computer Science, Electronics and Telecommunications of AGH University.

### v2

@@ IV-C Multilateration with Many Wires
A more robust algorithm requires a larger number of reference wires. In the following section, we present an approach feasible when we have at least six wires. Let us consider that all the wires are parallel to one of the axes of the coordinates system and call them , , or wires, according to the axis of their parallelism.
Now, let us consider a set of parallel wires, e.g., wires, like in Fig. 2. Again, the distances calculated in the ranging phase create cylinders of possible nano-machine positions along each wire. Thus, localization with Z wires enables to calculation of the (,) coordinates of the nano-machine with the following set of equations:
@@
, - coordinates of the nano-machine estimated with Z wires
- distance from the nano-machine to the wire
@@
Then, this matrix equation can be solved using the standard least-squares approach, obtaining the estimate vector containing the and coordinates of the nano-machine:
Now, the analogous approach can be performed for and wires. Having at least three wires and three wires, we can follow the algorithm given in (4)-(10) to obtain (,) and (,) coordinates. Each set of wires enables to calculation of two of the nano-machine coordinates only. However, assuming the number of , , and wires is respectively equal to , , and , the final nano-machine coordinates can be calculated as respective weighted means:
Here, it should be noted that it is sufficient to have just 6 wires in total, e.g, three wires, three wires and no wires. It means that is equal to 0 and the parameters and cannot be calculated, but according to (11) all three coordinates of the nano-machine can be estimated. Thus, the minimum number of wires for this version of the localization algorithm is equal to 6.
@@ V Simulation Results
To validate the proposed localization algorithms, a computer simulator is written in Python 3 programming language. The simulator reflects the geometric setting of DC wires being the magnetic sources (Sec. III-A) and nano-machines equipped with magnetometers operating inside of a human body (Sec. III-B). It also takes into account the Earth’s magnetic field which might be a source of measurement errors (Sec. V-A). For each position of the nano-machine inside of the body, the simulator calculates the measured magnetic field coming from each DC wire influenced by the Earth’s magnetic field. With these measurements, the position of the nano-machine is estimated based on the proposed localization algorithm (Sec. IV). Finally, the localization error is calculated in comparison to the real position of the nano-machine. Discussing it in more detail, each simulation follows the subsequent steps:
@@
Geometry definition. The general geometry of the simulation is created: the position of the DC wires and the relative inclination of the Earth’s magnetic field are established. It should be noted that the Biot-Savart law (1) is defined for infinitely long DC wires. In order to have accurate ranging estimations, the wires should be much longer than the human body, depending on the required accuracy even a few tens of meters long.
@@
Nano-machines positions and orientations. A voxel model of a 175 cm human male is implemented to emulate the human body. The model is depicted in Fig. 3; it is composed of 625,885 points in total, which results in the point spatial resolution of 0.5 cm along each axis. Each point represents a potential nano-machine position for which magnetic field measurements are performed and localization is attempted. At each point, a nano-machine is generated with a random orientation, which is repeated 100 times for statistical credibility.
@@
Localization and its error. The proposed localization algorithm is executed and the position of the nano-machine is estimated. The Euclidean distance between the real and estimated positions is calculated as the position error. It is averaged over 100 random nano-machine orientations at each point of the human body model.
V-A Earth’s Magnetic Field and Other Sources of Interference
@@
Ignoring the fact that Earth’s magnetic field exists and performing the localization would result in huge position estimation errors. We performed computer simulations to check this approach and the obtained position errors were about several dozen centimeters, which is at least an order of magnitude larger than the errors reported in Sec. V-B. Fortunately, a couple of models allow for the estimation of the geomagnetic field. In this paper, we use the commonly known World Mathematical Model (WMM) [60]. It was created by the U.S. National Oceanic and Atmospheric Administration’s National Centers for Environmental Information (NOAA/NCEI) and the British Geological Survey (BGS). The current model version was parameterized in 2020 and is supposed to be valid until 2025. The main goal of the WMM is to represent the planet’s magnetic field for all locations of the globe. What is quite important here, it also provides data on how large the model error is compared with the real Earth’s magnetic field. These errors for three orthogonal components (northern, eastern, and vertical ones) of the magnetic flux density are shown in Tab. LABEL:Table2 for the 2020 and 2025 years.
For the computer simulator, we calculate the average of the model error for 2020 and 2025 (see the third column of Tab. LABEL:Table2). Then, for each nano-machine, we randomly generate the WMM errors for all three components as uniformly distributed values from the range -error, +error. Thus, the final error of the measured magnetic field results from: (1) the WMM error, and (2) the relative measurement error of the magnetometers being 1% (see Sec. III-B).
Except for Earth’s magnetic field, some other sources can create interference for the magnetic field measured by magnetometers. The vast majority of magnetic field sources generate an alternating magnetic field; these are sources of electromagnetic waves or just devices powered with electrical AC current. It is assumed here that magnetometers are equipped with filters that pass a static magnetic field only. There are, however, some sources of static magnetic field, mostly cables with DC current. A low DC current is not going to create a significant magnetic field, as its field density decreases with distance. Nonetheless, some medical equipment like MRI devices can be supplied with a DC current even above 80 A, and thus placing the discussed localization system close to MRI devices, e.g., in the same room, should be clearly avoided.
@@ V-B Results
Several wire arrangements were investigated during simulations. The first one was the case with just 3 DC wires, as it is the minimal number of wires when the localization is feasible with the trilateration approach (Sec. IV-B). Then, scenarios with 6, 9, 15, and 30 DC wires are considered, for the more advanced algorithm, based on multilateration (Sec. IV-C). These wire arrangements create cage-like shapes where a patient can stand inside (see an example for 15 wires in Fig. 4) for medical investigation. The electrical current for all the wires was 100 A. The created magnetic flux density was never higher than 120 mT inside the human body, as it was the maximum value in the measurement range of the used magnetometers. Also, one scenario with a lower electrical current, only 10 A, was investigated; it resulted in a proportionally lower magnetic flux density inside the body. For each scenario and each of 625,885 nano-machine positions, 100 simulations were carried out and the average localization error was calculated. After performing Kolmogorov-Smirnov tests, it was noticed that, over different nano-machine positions, the localization errors did not have Gaussian distributions. Thus, their values are given with 1st, 2nd (median), and 3rd quartiles, together with maximum values. All these statistics are shown in Tab. LABEL:tab:loc_error.
It should be considered that nano-machines might move during the localization process. As it is assumed the nano-machines flow with the blood, their typical velocity is about 25 cm/s or lower, but in some rare cases can be even 50 cm/s. The time duration of the localization process mainly depends on the time of switching on and off the wires, as each wire should generate the static magnetic field separately. This issue was investigated performing additional calculations with COMSOL Multiphysics software. These investigations showed that even for the current of 100 A, the current stabilized after the time of about 1 . Thus, it might be assumed that the whole process of switching on, performing the measurements and switching off for even 30 wires is not going to be longer than 1 ms. Even for the rare occurring velocity of 50 cm/s, a nano-machine does not move further than 0.5 mm in that time period. This movement error is then treated as not significant, compared with localization errors reported below, being at least an order of magnitude larger.
In the 3 wires scenario, the median position error equals 5.01 cm, with the 1st and 3rd quartiles being 2.97 and 7.16 cm, respectively. The detailed map of the median error is depicted in Fig. 5. The error increases in the regions being far from the wires, especially the head, which is located about 170 cm from both and wires. It is due to the lower magnetic field there when generated by these wires, thus the disturbances resulting from the Earth’s field (see Sec. V-A) are more significant.
For the multilateration algorithm, the scenarios with 6, 9, 15, and 30 wires are compared in Tab. LABEL:tab:loc_error. The reported errors are much lower than for only 3 wires. Especially the scenario of 6 wires, which is the minimum number for multilateration, seems promising: the median position error is just 0.79 cm, with the 1st and 3rd quartiles being 0.7 and 0.95 cm, respectively. This proves the advantage of the multilateration approach over the simplistic trilateration. The error map for the 6 wires scenario is presented in Fig. 6. Here, the error is the highest in the feet area. It is because of the wires’ location: they are positioned around the body, but the feet are a body part farthest from the center of the wires’ system. In practical applications, it might be possible to adapt the size of the wires’ system to the height of the body decreasing the localization error. It might be also worthwhile to optimize the exact location of the wires. They should be positioned close to the body, to have a strong magnetic field there. Because of this reason, in this scenario, it was important to use a set of three wires, which are located along the body, and one of the other set of perpendicular ones, e.g. wires, as required for the localization algorithm. Probably, it is possible to obtain a slightly better localization accuracy after a careful optimization of the wires’ positions. This might be done e.g. with gradient methods, but it depends on the exact shape of the body and it is out of the scope of this study.
Further increasing the number of wires improves the localization accuracy only slightly. For 9 wires, the median error (2nd quartile) is 0.77 and it remains the same, accurate to two decimal places, for 15 wires. In the scenario with 30 wires, the median error decreases to 0.69 cm, so there is an improvement, but it is just a 13% lower error compared with 6 wires.
Finally, a case of a lower DC current was investigated. Two scenarios were compared for 15 wires: apart from the current of 100 A, a case of 10 A was simulated as well (see Tab. LABEL:tab:loc_error). For the lower current in the wires, the localization error significantly increases. The median position error grows from 0.77 cm (100 A) to 3.05 cm (10 A). This situation can be explained by the disturbances created by the Earth’s magnetic field. The lower DC current in the wires results in a lower magnetic field density, thus the error of the WMM (Sec. V-A) model is relatively stronger.
| Scenario | 3 wires, 100 A | 6 wires, 100 A | 9 wires, 100 A | 15 wires, 100 A | 15 wires, 10 A | 30 wires, 100 A |
| Errors [cm]: | ||||||
| Maximum X error | 10.88 | 0.97 | 1.00 | 0.89 | 4.80 | 0.76 |
| Maximum Y error | 7.63 | 0.50 | 0.92 | 0.94 | 2.40 | 0.93 |
| Maximum Z error | 9.78 | 1.88 | 1.55 | 1.29 | 10.64 | 1.25 |
| Maximum position error | 16.48 | 2.18 | 2.15 | 1.78 | 11.11 | 1.75 |
| Median X error | 2.12 | 0.38 | 0.38 | 0.37 | 0.83 | 0.31 |
| Median Y error | 1.71 | 0.24 | 0.38 | 0.39 | 0.87 | 0.36 |
| Median Z error | 3.48 | 0.56 | 0.42 | 0.36 | 2.46 | 0.33 |
| Median position error | 5.01 | 0.79 | 0.77 | 0.77 | 3.05 | 0.69 |
| 1st quartile X error | 1.40 | 0.33 | 0.34 | 0.30 | 0.53 | 0.24 |
| 1st quartile Y error | 1.04 | 0.22 | 0.33 | 0.33 | 0.61 | 0.29 |
| 1st quartile Z error | 2.07 | 0.46 | 0.34 | 0.26 | 1.18 | 0.22 |
| 1st quartile position error | 2.97 | 0.70 | 0.68 | 0.60 | 1.82 | 0.52 |
| 3rd quartile X error | 3.23 | 0.45 | 0.45 | 0.46 | 1.28 | 0.40 |
| 3rd quartile Y error | 2.82 | 0.26 | 0.44 | 0.48 | 1.12 | 0.47 |
| 3rd quartile Z error | 4.91 | 0.69 | 0.55 | 0.49 | 3.84 | 0.47 |
| 3rd quartile position error | 7.16 | 0.95 | 0.92 | 0.94 | 4.33 | 0.87 |
@@ VI Conclusions
In this paper, we presented a new magnetic-field-based localization system suitable for in-body nano-machines. In the system, external electric wires with constant current generate a magnetic field that can be measured by tiny magnetometers mounted on board the nano-machines. These measurements are sent out of the body through a body-area network and are used to calculate the nano-machine positions. We proposed two versions of the localization algorithm: for three, and six or more electric wires. We also explained how to take into account Earth’s magnetic field. We performed computer simulations for both algorithm versions showing that even with errors induced by the magnetometers and Earth’s magnetic field, the obtained localization accuracy can be about 1 cm or better. The proposed magnetic localization approach is a solution that fits well with the currently increasing number of in-body medical systems. Recently, micro- and nano-scale robots for drug delivery are gaining attention, together with techniques where the magnetic field is used for their control and locomotion [61]. It can also observe the development of systems based on ingestible electronics proposing an opportunity to monitor gastrointestinal diseases in real time [62]. For all these devices, the presented localization system can offer their precise position estimation, expanding their capabilities and applications in personalized medical treatment.
Future works extending these studies may go in many different directions ranging from specific system applications through nano-device hardware fine-tuning to new networking architectures. One particularly interesting example is substituting the external DC wires, generating the magnetic field, with in-body magnets. Such magnets are already considered for guiding some nano-particles [63]. The advantage of this solution is the fact that no external reference wires are required, as the source of the magnetic field is inside the body. It is, however, still to be investigated if the magnets’ positions could be sufficiently stable and the fields suitably regular and strong for accurate localization. Another related research line is a study on generating a uniform magnetic field around external wires of finite length. Here we must assume the wires are quite long, even a few tens of meters, but for practical reasons it would be more convenient to use shorter wires. Thus, it would be worth examining what is the localization error resulting from the non-uniform magnetic field and how a uniform field can be generated in a more compact wires’ set-up. When considering a specific number of wires and a particular human body, the wires’ exact positions could be optimized, e.g. with gradient methods, for the most uniform magnetic field and the best localization performance (moving the wires closer to the body center and adapting their positions to the body shape). These examples of future studies just indicate a wide range of research topics related to magnetic field applications for in-body medical diagnosis.
@@ Acknowledgement
The authors would like to thank Andrzej Kozłowski for his invaluable comments and suggestions regarding magnetometers.